# Monodromy Defects in Maximally Supersymmetric Yang-Mills Theories from Holography

### v1

@@ 4 The $p=5$ case
The solution of [34] is obtained in the 7D U(1)2 gauged supergravity. In this case, theory contains the metric, two U(1) gauge fields ($A^{I}$), and two scalars ($\lambda^{I}$), with $I=1,2$. The background configuration can be then written as
@@ 4.1 Vacuum Solution
comparing with (4.1b) we identify $g=1/\ell$. This condition do not change for the defect solution. Notice that, in brane frame, the $g_{zz}$ component does not depend on $y$, we comment on this feature below. The vacuum is parametrized in a similar way to (2.14a)-(2.14c), where the RG flow information is encoded both in $(y,\rho)$ via (2.13). In this case, it is also possible to move the vacuum to the canonical frame of via
@@ 4.2 Boundary Conditions
We now proceed to the study of boundary conditions for solution. At $y=y_{\text{core}}$ these are the same as in 3.1, while for $y\rightarrow+\infty$, the condition for the gauge field remains the same, but the metric must asymptote to the vacuum studied in the section above.
@@
In this way, the parameters characterizing the solution are $p$, $n$, $y_{\text{core}}$ and $r_{1}$, with $y_{\text{core}}>0$ and $r_{1}<0$. Of course, one could instead invert the relation between the parameters above to $(y_{\text{core}},r_{1})\rightarrow\mu^{I}$, to have the later ones as free parameters, but we find this unnecessary complicated.
@@ 5 Defect Entanglement Entropy
Here, we finish our study by applying the free energy formula [^14] This formula was proposed as a shortcut to the free energy computation via entanglement entropy. of [35, 36] to the codimension-2 monodromy defect solutions of Section 3. This requires choosing a holographic direction that realizes the holographic RG flow. For defect backgrounds studied here, which asymptote to the vacuum (2.14a)-(2.14c), the RG-flow information is encoded in the coordinates $(\rho,y)$, so that a direct computation of the central charge/ free energy of the $(p+1)$-dimensional theory with a defect is not possible in this parametrization.
@@ 5.1 Renormalization Scheme
We apply this formula for the solutions studied in Section 3, interpreting them solutions as duals to a (p1)(p-1)-dimensional QFT, hence treating $y$ and $z$ as part of the internal manifold. Using the uplifts in Appendix A, for $p=2,3,4$, we find
@@
where we denote by $c^{(p)}_{\text{hol}}$ the free energy/Weyl anomaly of the $(p+1)$-dimensional theory, which was obtained in (2.10).
@@ 5.2 Supersymmetric Surface Defect in 4D SYM
from where it can be showed that the following relation holds
@@ 5.4 Supersymmetric Line Defect in 3D SYM
$c^{\text{renorm}}_{\text{hol}}$ $=C_{2}\frac{1}{224\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}n^{2}}\bigg((g\mu_{F_{1}}-g\mu_{F_{2}}-2)(2g\mu_{F_{1}}+g\mu_{F_{2}}+2)(g\mu_{F_{1}}+2g\mu_{F_{2}}-2)-232n^{3}$ (5.29)
$-456n^{2}+54n\left(g\mu_{F_{1}}^{2}+g\mu_{F_{1}}g\mu_{F_{2}}+g\mu_{F_{2}}^{2}-4\right)+224n^{2}\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}\bigg)c_{\text{hol}}^{(2)},$
@@
$\Delta_{1}=2(1+n)-g\mu_{F_{1}}-2g\mu_{F_{2}},\qquad\Delta_{2}=2(1+n)-g\mu_{F_{1}}+g\mu_{F_{2}},\qquad\Delta_{3}=2(1+n)+g\mu_{F_{1}}+2\mu_{F_{2}}.$ (5.30)
@@ 6.1 On the non-conformal dEE
It is important to remark that this process does not fit into the description of [49]: at the level of the lower dimensional gauged supergravity, both are 4-dimensional, the difference being the field content and the gauge group. Still, it would be interesting if a relation similar to (5.22) and (6.1) holds to the codimension-2 defect in the D2-brane theory, that is
@@
${\cal{I}}^{(1)}$ $=\kappa_{d}\frac{49\pi}{96n}\left(16n^{2}-\Delta_{1}^{2/3}\Delta_{2}^{2/3}\Delta_{3}^{2/3}\right)c^{(2)}_{\text{hol}}$,
$h_{D}^{(1)}$ $=\kappa_{d}\frac{35}{128n^{2}}\frac{1}{\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}}\big(8(n+1)n^{2}-(g\mu_{F_{1}}-g\mu_{F_{2}}-2)(2g\mu_{F_{1}}+g\mu_{F_{2}}+2)(g\mu_{F_{1}}+2g\mu_{F_{2}}-2)$
$+2n\left(g^{2}\mu_{F_{1}}^{2}+g^{2}\mu_{F_{1}}\mu_{F_{2}}+g^{2}\mu_{F_{2}}^{2}-4\right)\big)c^{(2)}_{\text{hol}}.$ (6.6)
@@ A.1 Type II Supergravity
where $(F_{\pm},F_{\pm})_{8}$ is the 8-form part of $F_{\pm}\wedge\lambda(F_{\pm})$, $d_{H}=d+H\wedge$. We have defined for a $k$-form $X$
$(X_{k})_{M}:=\iota_{dx^{M}}X_{k},~~~~X_{k}^{2}:=\sum_{k}\frac{1}{k!}(X_{k})_{M_{1}\dots M_{k}}(X_{k})^{M_{1}\dots M_{k}},$ (A.4)
@@
$\left(\nabla-\frac{1}{4}H-d\Phi\right)\epsilon_{1}=0,\quad\left(\nabla+\frac{1}{4}H-d\Phi\right)\epsilon_{2}=0.$ (A.6b)
@@
The solutions in Section 3.4 are obtained using a U(1)2 truncation of the maximally supersymmetric 6D SO(5) gauged supergravity [68]. The bosonic content of this truncation is: the metric, two gauge fields $A^{I}$, and three real scalars $\lambda_{I}$ and $\sigma$, where $I=1,2$. The bosonic lagrangian reads [34]
@@
Finally, $ds^{2}_{6}$ is the 6-dimensional gauged supergravity metric and vol6 is its volume form.
@@
while $ds^{2}_{5}$ and vol5 are correspond to the 5-dimensional gauged supergravity solution.
@@
We now present the conventions for the solution in Section 3.2 that can be see as the $U(1)^{3}$ truncation of the 4d ISO(7) gauged theory, which can be obtained as a deformation of the maximal SO(8) gauging [40]. The bosonic fields are the metric, three gauge fields $A^{I}$ and four real scalars $\phi_{i}$. The bosonic action is [34]
@@
$\mathcal{V}=-\frac{g^{2}}{2}\left[\sum_{0\leq i<j\leq 3}X^{i}X^{j}-\frac{1}{2}(X^{0})^{2}\right].$ (A.28)

### v2

@@ 4 The $p=5$ case
The solution of [15] is obtained in the 7D U(1)2 gauged supergravity. In this case, the theory contains the metric, two U(1) gauge fields ($A^{I}$), and two scalars ($\lambda^{I}$), with $I=1,2$. The background configuration can be then written as
@@ 4.1 Vacuum Solution
comparing with (4.1b) we identify $g=1/\ell$. This condition do not change for the defect solution. Notice that, in brane frame, the $g_{zz}$ component does not depend on $y$, we comment on this feature below. The vacuum is parametrized in a similar way to (2.14a)-(2.14c), where the RG flow information is encoded both in $(y,\rho)$ via (2.13). In this case, it is also possible to move the vacuum to the canonical frame via
@@ 4.2 Boundary Conditions
We now proceed to the study of boundary conditions for this solution. At $y=y_{\text{core}}$ these are the same as in 3.1, while for $y\rightarrow+\infty$, the condition for the gauge field remains the same, but the metric must asymptote to the vacuum studied in the section above.
@@
In this way, the parameters characterizing the solution are $p$, $n$, $y_{\text{core}}$ and $r_{1}$, with $y_{\text{core}}>0$ and $r_{1}<0$. Of course, one could instead invert the relation between the parameters above to $(y_{\text{core}},r_{1})\rightarrow\mu^{I}$, to have the later ones as free parameters, but we find this unnecessarily complicated.
@@ 5 Defect Entanglement Entropy
Here, we finish our study by applying the free energy formula [^15] This formula was proposed as a shortcut to the free energy computation via entanglement entropy. of [62, 4] to the codimension-2 monodromy defect solutions of Section 3. This requires choosing a holographic direction that realizes the holographic RG flow. For the defect backgrounds studied here, which asymptote to the vacuum (2.14a)-(2.14c), the RG-flow information is encoded in the coordinates $(\rho,y)$, so that a direct computation of the central charge/ free energy of the $(p+1)$-dimensional theory with a defect is not possible in this parametrization.
@@ 5.1 Renormalization Scheme
We apply this formula for the solutions studied in Section 3, interpreting them as duals to a (p1)(p-1)-dimensional QFT, hence treating $y$ and $z$ as part of the internal manifold. Using the uplifts in Appendix A, for $p=2,3,4$, we find
@@
where we used the free energy of the $(p+1)$-dimensional theory, computed in (2.10).
@@ 5.2 Supersymmetric Surface Defect in 4D SYM
from where it can be shown that the following relation holds
@@ 5.4 Supersymmetric Line Defect in 3D SYM
$c^{\text{renorm}}_{\text{hol}}$ $=C_{2}\frac{1}{28\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}n^{2}}\bigg((g\mu_{F_{1}}-g\mu_{F_{2}}-1)(2g\mu_{F_{1}}+g\mu_{F_{2}}+1)(g\mu_{F_{1}}+2g\mu_{F_{2}}-1)-29n^{3}$ (5.29)
$-57n^{2}+27n\left((g\mu_{F_{1}})^{2}+g\mu_{F_{1}}g\mu_{F_{2}}+(g\mu_{F_{2}})^{2}-1\right)+56n^{2}\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}\bigg)c_{\text{hol}}^{(2)},$
@@
$\Delta_{1}=n+1-g\mu_{F_{1}}-2g\mu_{F_{2}},\qquad\Delta_{2}=n+1-g\mu_{F_{1}}+g\mu_{F_{2}},\qquad\Delta_{3}=n+1+g\mu_{F_{1}}+2\mu_{F_{2}}.$ (5.30)
@@ 6.1 On the non-conformal dEE
It is important to remark that this process does not fit into the description of [43]: at the level of the lower dimensional gauged supergravity, both are 4-dimensional, the difference being the field content and the gauge group. Still, it would be interesting if a relation similar to (5.22) and (6.1) holds for the codimension-2 defect in the D2-brane theory, that is
@@
${\cal{I}}^{(1)}$ $=\kappa_{d}\frac{49\pi}{24n}\left(4n^{2}-\Delta_{1}^{2/3}\Delta_{2}^{2/3}\Delta_{3}^{2/3}\right)c^{(2)}_{\text{hol}}$,
$h_{D}^{(1)}$ $=\kappa_{d}\frac{35}{8n^{2}}\frac{2}{\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}}\big((n+1)n^{2}-(g\mu_{F_{1}}-g\mu_{F_{2}}-1)(2g\mu_{F_{1}}+g\mu_{F_{2}}+1)(g\mu_{F_{1}}+2g\mu_{F_{2}}-1)$
$+n\left((g\mu_{F_{1}})^{2}+g\mu_{F_{1}}g\mu_{F_{2}}+(g\mu_{F_{2}})^{2}-1\right)\big)c^{(2)}_{\text{hol}}.$ (6.6)
@@ A.1 Type II Supergravity
where $(F_{\pm},F_{\pm})_{8}$ is the 8-form part of $F_{\pm}\wedge\lambda(F_{\pm})$, $\mathrm{d}_{H}=\mathrm{d}+H\wedge$. We have defined for a $k$-form $X$
$(X_{k})_{M}:=\iota_{\mathrm{d}x^{M}}X_{k},~~~~X_{k}^{2}:=\sum_{k}\frac{1}{k!}(X_{k})_{M_{1}\dots M_{k}}(X_{k})^{M_{1}\dots M_{k}},$ (A.4)
@@
$\left(\nabla-\frac{1}{4}H-\mathrm{d}\Phi\right)\epsilon_{1}=0,\quad\left(\nabla+\frac{1}{4}H-\mathrm{d}\Phi\right)\epsilon_{2}=0.$ (A.6b)
@@
The solutions in Section 3.4 are obtained using a U(1)2 truncation of the 6D SO(5) gauged supergravity [30]. The bosonic content of this truncation is: the metric, two gauge fields $A^{I}$, and three real scalars $\lambda_{I}$ and $\sigma$, where $I=1,2$. The bosonic lagrangian reads [15]
@@
Finally, $\mathrm{d}s^{2}_{6}$ is the 6-dimensional gauged supergravity metric and vol6 is its volume form.
@@
while $\mathrm{d}s^{2}_{5}$ and vol5 are correspond to the 5-dimensional gauged supergravity solution.
@@
We now present the conventions for the solution in Section 3.2 that can be see as the U(1)3 truncation of the 4D ISO(7) gauged theory, which can be obtained as a deformation of the maximal SO(8) gauging [53]. The bosonic fields are the metric, three gauge fields $A^{I}$ and four real scalars $\phi_{i}$. The bosonic action is [^24] We use conventions such that the gauge coupling $g_{\text{here}}=\frac{1}{2}g_{\text{there}}$. [15]
@@
$\mathcal{V}=-2g^{2}\left[\sum_{0\leq i<j\leq 3}X^{i}X^{j}-\frac{1}{2}(X^{0})^{2}\right].$ (A.28)